# ImageSet2Text: Describing Sets of Images through Text

### v1

@@ Abstract
We introduce ImageSet2Text, a novel approach that leverages vision-language foundation models to automatically create natural language descriptions of image sets. Inspired by concept bottleneck models (CBMs) and based on visual-question answering (VQA) chains, ImageSet2Text iteratively extracts key concepts from image subsets, encodes them into a structured graph, and refines insights using an external knowledge graph and CLIP-based validation. This iterative process enhances interpretability and enables accurate and detailed set-level summarization. Through extensive experiments, we evaluate ImageSet2Text´s descriptions on accuracy, completeness, readability and overall quality, benchmarking it against existing vision-language models and introducing new datasets for large-scale group image captioning.
Figure 1: ImageSet2Text generates detailed and nuanced descriptions from large sets of images. We report two descriptions and corresponding 4x4 grids, composed of 16 randomly selected images belonging to the described image sets [55] [57].
@@ 1 Introduction
Summarizing entire image sets in natural language is important to capture overarching themes and trends among the images, hence simplifying the navigation and understanding of large image collections. In fact, image set summarization is necessary in a variety of applications, including assistive technologies [5, 22] and cultural analytics [7, 40]. In explainable AI, dataset-level insights have been found to be valuable for bias detection, influential sample analysis, and data segmentation [47, 54, 12, 15, 19]. Furthermore, the growing demand for transparency in AI training datasets has made this need even more pressing [20], especially with emerging AI regulations such as the EU AI Act [48].
While significant progress has been made in vision-language modeling and image captioning, most approaches focus on individual images [25, 61, 64] or sets with a limited number of images [10, 2, 30, 65]. Despite its importance, the summarization of large image sets remains an unsolved problem due to fundamental technical limitations: existing methods are typically not designed to efficiently process multiple visual inputs at once, limiting their ability to extract holistic insights from image collections [18].
In this paper, we propose ImageSet2Text, a novel approach that leverages vision-language foundation models to generate natural language descriptions [^1] We refer to captions as “short pieces of text” [6], while descriptions are typically longer and more detailed. of large-scale image sets. ImageSet2Text is inspired by the rationale behind concept bottleneck models (CBMs) [29], which are typically used for classification tasks. CBMs predict intermediate concepts before making a final decision, ensuring that information flows through a human-interpretable representation [52]. While traditional CBMs are based on predefined sets of concepts, recent research has explored more flexible formulations, relying on visual-question answering (VQA) chains [56, 9].
Similarly, ImageSet2Text leverages the strong capability to align visual and textual representations of pretrained multimodal foundation models [51, 2] to generate a comprehensive description of an image set over multiple iterations. In each cycle, a random subset of images is selected to perform Large Language Model (LLM)-based VQA. The extracted information is then encoded into a graph that contains the key concepts from the answers. ImageSet2Text also integrates an external knowledge graph [44] to hypothesize relevant information and validates these hypotheses on the entire image set using contrastive vision-language (CVL) embeddings [51]. The information that is confirmed on a large portion of the images is added to the graph and used to seed the next VQA iteration. By combining VQA chains, graph-based concept representations, and iterative refinement, ImageSet2Text enhances the interpretability of large image sets and enables novel applications in image set understanding.
Through extensive experiments, we evaluate ImageSet2Text´s descriptions according to their accuracy, completeness, readability and overall quality. To measure accuracy, we propose two datasets for large-scale group image captioning, and we benchmark ImageSet2Text on these datasets against existing vision-language models. We assess completeness by means of an image set comparison task. Finally, we perform a user study with 200 participants to collect human feedback on the readability and overall quality of ImageSet2Text´s descriptions.
@@ 2 Related Work
Image Captioning aims to generate semantically meaningful, short textual descriptions of images by recognizing objects, scene types, object properties, and relationships [25] in the images. Standard image captioning methods are mainly based on feature learning through deep learning [61, 64]. Recently, Zhu et al. [68] introduced ChatCaptioner, combining VQA with chat logs to iteratively refine captions. In addition, Mao et al. [41] explored context-aware captioning, proposing to generate captions tailored to user-defined contexts. The role of context has been explored particularly in Art History captioning research, where authors have analyzed how captions can vary based on interpretations [4, 7, 39].
@@
Group-Image Captioning extends single-image captioning to small collections of images (typically from 2 to 30 images), with the aim of identifying and summarizing similarities among them [10, 2, 30, 65]. Different approaches have been proposed in the literature, including incorporating a temporal relationship among images [63] and understanding the difference between image pairs [8, 28, 46] or target and reference image groups [32]. Scene graph representations have also been employed to model the relationships between elements in the images and summarize such relationships among more images [50, 49]. Recent work has also shown the potential of LLMs to perform individual and small-group image captioning [1]. In parallel to these efforts, the evaluation benchmarks for group image captioning focus on spatial, semantic, and temporal aspects related to small groups of images [43] or on evaluating large vision-language models on multi-image question answering [38]. Despite this variety of approaches, the main limitation of current methods is their ability to handle groups with larger numbers (hundreds to thousands) of images, leaving the summarization of such image sets an open challenge [49].
Understanding Collections of Images is crucial in an era of large-scale visual data, but efforts in this direction are still limited. Research on describing large image sets has focused primarily on concept-level prototypes [17, 59], color-based statistical analysis [58], and set-level classification [62]. However, these approaches do not generate easy-to-interpret textual descriptions. A step towards bridging this gap was taken by Dunlap et al., who introduced the new task of Set Difference Captioning [18] (SDC). This new task consists of comparing two image sets and generating a caption that applies to one of the sets but not the other. In our paper, we contribute to this field by proposing a novel method to generate textual descriptions of image collections with hundreds to thousands of examples per group, moving beyond comparative descriptions and towards comprehensive set-level insights.
@@
Foundation Models are increasingly used to solve complex vision-language tasks. In addition to [18], querying a VQA model through an LLM has been used to iteratively improve image and video captions [11, 68], to propose an open set bias detection technique in text-to-image generation [16] and to evaluate text-to-image generation faithfulness [26]. ImageSet2Text aligns with these methods by integrating multiple foundation models to generate textual descriptions from image collections, leveraging both vision-language reasoning and iterative refinement mechanisms.
@@ 3 ImageSet2Text
ImageSet2Text aims to generate nuanced natural language descriptions of image sets that highlight the common visual elements present in most of the images. As shown in Fig. 2, ImageSet2Text leverages structured prompts in LLMs, VQA, knowledge graphs, and CVL models.
Formally, given a set of $N$ images $\mathcal{D}=\{x_{1},\dots,x_{N}\}$, ImageSet2Text automatically generates a textual description $d$ that summarizes the visual elements in $\mathcal{D}$. This is achieved by constructing an intermediate graph represented as a list of triplets $\mathcal{G}=\{\langle s,p,o\rangle_{1},...,\langle s,p,o\rangle_{T}\}$, where each triplet consists of a subject $s$, a predicate $p$, and an object $o$ that capture the key visual elements and their interactions in $\mathcal{D}$. To build $\mathcal{G}$, ImageSet2Text follows an iterative process with $T$ iterations depicted in Fig. 2, with the following steps:
@@
    Initialization ($\tau=0$): The initial $\mathcal{G}_{0}$ contains a root node, $s_{0}=\text{`{image}'}$ ’, linked to three pending predicates, $p_{0}=\text{`{content}'}$ ’, $p_{1}=\text{`{background}'}$ ’, and $p_{2}=\text{`{style}'}$ ’.
@@
    Iterations ($\tau=1,...,T-1$): Each iteration is composed of two phases:
@@
      Guess what is in the set – A random subset of images $\mathcal{S}\subset\mathcal{D}$, with $|\mathcal{S}|=M\ll N$, is analyzed to hypothesize elements present in $\mathcal{D}$, where $M$ is a predefined parameter.
@@
      Look and keep – The formulated hypothesis is validated on the entire set $\mathcal{D}$. If confirmed, it is used to update $\mathcal{G}_{\tau}$.
@@
    Termination ($\tau=T$): After convergence at $\tau=T$, the final graph representation $\mathcal{G}=\mathcal{G}_{T}$ is obtained. Finally, a coherent and concise textual description $d$ is generated from $\mathcal{G}$.
@@
Figure 2: Overview of ImageSet2Text, considering an example set from the PairedImageSets datasets [18]. The figure shows how the different modules of the iterative process allow inferring information from the input image set, eventually generating a nuanced textual description.
@@ 3.1 Guess what is in the Set
The first phase generates a set of hypotheses from the random sample $\mathcal{S}$ of $M$ images for later validation on the full set $\mathcal{D}$. Let $\tau$ be the current time step. Given the current graph $\mathcal{G}_{\tau}$, ImageSet2Text selects the closest leaf node to the root node as the predicate $p$ for analysis at this step, along with its parent node $s$ as the corresponding subject.
@@
An LLM is prompted to ask a specific question about the images in $\mathcal{S}$ depending on the values of $s$ and $p$. We provide an illustrative example in Fig. 2. Consider the set $\mathcal{S}$ of images of a desert in a certain iteration. With $s=\text{`{light}'}$ ’ and $p=\text{`{color}'}$ ’, the formulated question is:
@@
The question is applied to all the images $x_{i}\in\mathcal{S}$, resulting in a set of answers $\mathcal{A}=\{a_{1},a_{2},\dots,a_{M}\}$, where $a_{i}$ denotes the answer for $x_{i}$.
@@
Considering the answers in $\mathcal{A}$, the goal of each iteration is to verify whether the predicate under consideration $p$ can expand $\mathcal{G}_{\tau}$. To achieve this, the LLM is prompted to perform two tasks: 1) summarization, i.e., condense $\mathcal{A}$ into a single hypothesis $h$, which has to be formulated as a triplet where the subject $s$ and predicate $p$ are given, while the object $o$ has to be derived from $\mathcal{A}$; and 2) completion, i.e., suggest possible continuations for the hypothesis $h$ in the form of a list of predicates $P$, defining potential expansions of $\mathcal{G}_{\tau}$ from $o$. Following the previous example, the set of answers provided to the question would yield hypothesis $h$ and continuations $P$:
$h=\langle\text{`{desert}'},\text{`{color}'},\text{`{golden yellow}'}\rangle$ ’ ⟩ $P=\{\text{`{shade}'},\text{`{brightness}'}\}$ ’ }
@@
Since $h$ is derived from the subset $\mathcal{S}$, it may not generalize well to the full image set $\mathcal{D}$ due to sampling bias. To mitigate this, ImageSet2Text creates a set =$\mathcal{H}=\{h_{0},h_{1},\dots,h_{k}\}$ of hypotheses ordered from more general ($h_{k}$) to more specific ($h_{0}$), where $h_{k}\supset h_{k-1}\supset\dots\supset h_{0}$ and $h_{0}=h$. To generate $\mathcal{H}$, ImageSet2Text relies on a lexical knowledge graph $KG$. Let $KG=(V,R)$ be a directed graph, where $V$ is the set of lexical entries and $R$ represents semantic relations between nodes in $V$. For any given node $v\in V$, its parent node represents a more general concept (hypernym), while its children nodes represent more specific concepts (hyponyms). In addition, two nodes $v_{1},v_{2}\in V$ are sibling nodes if they share the same parent node, but correspond to different lexical concepts.
The set $\mathcal{H}$ is obtained by traversing upward in the knowledge hierarchy of $KG$ by a maximum number of steps $\delta$. Each hypothesis $h_{i}$ is derived by generalizing the object $o$ based on the hypernyms in $KG$. In sum, given a hypothesis $h_{i}=\langle s,p,o_{i}\rangle$, its generalization $h_{i+1}$ is created as $h_{i+1}=\langle s,p,o_{i+1}\rangle$, where $o_{i}=\text{parent}^{i}(o)$ and the parent function is the operation of moving to the hypernym of a lexical entry in $KG$.
In the ongoing example, the hypotheses in $\mathcal{H}$ follow the hierarchy:
$h_{0}=\langle\text{`{desert}'},\text{`{color}'},\text{`{golden yellow}'}\rangle$ ’ ⟩ $h_{1}=\langle\text{`{desert}'},\text{`{color}'},\text{`{warm color}'}\rangle$ ’ ⟩

### v2

@@ Abstract
In the era of large-scale visual data, understanding collections of images is a challenging yet important task. To this end, we introduce ImageSet2Text, a novel method to automatically generate natural language descriptions of image sets. Based on large language models, visual-question answering chains, an external lexical graph, and CLIP-based verification, ImageSet2Text iteratively extracts key concepts from image subsets and organizes them into a structured concept graph. We conduct extensive experiments evaluating the quality of the generated descriptions in terms of accuracy, completeness, and user satisfaction. We also examine the method’s behavior through ablation studies, scalability assessments, and failure analyses. Results demonstrate that ImageSet2Text combines data-driven AI and symbolic representations to reliably summarize large image collections for a wide range of applications.
Figure 1: ImageSet2Text generates detailed and nuanced descriptions from large sets of images. We report exemplary generated descriptions for two groups of images (sharma2018; artgan2018).
11footnotetext: Equal contribution. Work conducted at ELLIS Alicante.
Codehttps://github.com/ellisalicante/ImageSet2Text
Datasetscode+/tree/main/data
Extended versionhttps://arxiv.org/pdf/2503.19361
@@ 1 Introduction
The analysis of large-scale image sets is essential to uncover visual patterns that isolated samples fail to reveal (deng2025visual). For example, describing a dataset of historical newspaper photographs can disclose stylistic trends over time or systematic biases (such as underrepresented demographics or recurring stereotypical depictions), which only become apparent when analyzing the images collectively. To make such analyses scalable to large image collections, automated tools are indispensable. While visualizations help identify broad patterns (manovich2012), automatically generated textual summaries can transform overwhelming visual collections into interpretable knowledge.
A variety of application domains would benefit from automatic summarization of image collections, including assistive technologies (bigham2010; gurari2020), cultural analytics (shen2019discovering; cetinic21; manovich2020), bias detection (khosla2012undoing), socio-economics (jean2016combining; dubey2016deep; deng2025visual), exploratory data analysis (boiman2007detecting), and training data transparency (gebru2021) — whose need is intensified by emerging AI regulations (EU_AI_Act_2024). In addition, in explainable AI, dataset-level insights have been found to be valuable for influential sample analysis and data segmentation (park2023; shah2023; chung2019; d2022; eyuboglu2022).
However, while large vision-language models have made significant progress in the last few years (achiam2023; bai:2025), their application to describing large-scale image collections remains limited. Current approaches typically handle either individual images for captioning (hossain2019; vinyals2015; xu2015) or descriptions of small curated sets (Chen_2018; alayrac2022; li2023mimic; yao2022), failing to address the challenges of summarizing large visual datasets. This limitation remains unresolved due to fundamental technical challenges in processing multiple visual inputs at once (visdiff; deng2025visual).
In this paper, we propose ImageSet2Text, a method for generating natural language descriptions [^1] We refer to captions as “short pieces of text” (cambridge), while descriptions are longer and more detailed. (pi2024) of large collection of images, as shown in Fig. 1. ImageSet2Text leverages multimodal large language models (LLMs) through an iterative visual question answering (VQA) process, which combines hypothesis formulation-verification with external knowledge to extract comprehensive insights from image collections. Inspired by recent concept bottleneck models (CBMs) (koh2020; tan2024; chattopadhyay2024), ImageSet2Text extends this interpretable approach based on intermediate open-set concept prediction beyond classification tasks.
The descriptions generated by ImageSet2Text are obtained by means of an iterative process, consisting of two phases: (a) Guess what is in the set, and (b) Look and keep. To ensure scalability, a small subset of images is randomly selected in each iteration. In Guess what is in the set, an LLM-based VQA process identifies prevalent visual elements within the subset, and an external lexical graph is integrated to formulate hypotheses about the original set of images. In Look and keep, the hypotheses are verified using contrastive vision-language (CVL) embeddings (radford2021) to assess their consistency across the entire image set. Verified hypotheses are joined into a concept graph and used to seed the next iteration. The process terminates when the concept graph cannot be updated anymore, at which point a final description of the entire image set is generated based on the accumulated information.
In extensive experiments (see Fig. 2), we evaluate ImageSet2Text´s descriptions according to their: 1) accuracy, via a large-scale group image captioning experiment; 2) completeness, by means of an image sets comparison task; and 3) user satisfaction, through a user study. We also evaluate its behavior via: 4) an ablation study, 5) a scalability estimate, and 6) an analysis of failure cases. Our results indicate that the generated descriptions accurately capture the visual content of large collections, offering rich detail and human-friendly readability, highlighting the potential of ImageSet2Text for diverse applications.
Figure 2: ImageSet2Text’s evaluation considers three key properties of the descriptions (accuracy, completeness and user satisfaction) and analyzes the method’s behavior through ablations, scalability estimates and failure cases analysis, showing great versatility for potential applications.
@@ 2 Related Work
Image Captioning
generates short textual descriptions of images by recognizing objects, attributes, scenes, and their relationships (hossain2019). Early methods relied on deep learning for feature extraction (vinyals2015; xu2015), while recent approaches leverage LLMs. For example, ChatCaptioner integrates VQA with chat logs for iteratively refinement (zhu2023), and mao2024 (mao2024) proposes the creation of context-aware, user-specific captions. The importance of context has been highlighted in image captioning for Art History, where different interpretations can lead to different descriptions (bai21; cetinic21; yue24).
Group-Image Captioning
@@
extends single-image captioning to small sets (typically 2302-30 images) by identifying shared patterns in the images (Chen_2018; alayrac2022; li2023mimic; yao2022). Proposed methods include modeling temporal relationships among images (wang2019), analyzing pairwise differences (chang2023; kim2021; park2019), and comparing target and reference groups of images (li2020). Scene graphs have also been used to model and summarize the relationships between visual elements (phueaksri2024; phueaksri2023), while LLMs appear to be promising for both individual and small-group captioning (achiam2023). Benchmarks focus on spatial, semantic, and temporal aspects of small groups of images (meng2024) or on evaluating large vision-language models in multi-image question answering (liu2024). Despite the variety of approaches, the main limitation remains scaling to larger groups containing hundreds or thousands of images (phueaksri2023).
Understanding Collections of Images
remains challenging despite its importance in today’s world of large-scale visual data. Existing approaches, such as textual PCA (hupert2022), concept-level prototypes (doersch2015; van2023), color-based statistical analysis (torralba2011), and set-level classification (wang2022) fail to produce interpretable textual descriptions. Steps towards bridging this gap include set difference captioning (SDC) (visdiff) and temporal change detection in urbanist images (deng2025visual). In this paper, we advance beyond comparative approaches by proposing a novel method for generating comprehensive textual descriptions of large image collections.
Foundation Models
@@
are increasingly used for complex vision-language tasks. In addition to SDC (visdiff), querying a VQA model through an LLM has been used to iteratively improve image and video captions (chen2023; zhu2023), to detect biases in text-to-image generation (d2024), and to evaluate text-to-image generation faithfulness (hu2023). ImageSet2Text extends this paradigm by integrating multiple foundation models to describe image collections, contributing to emerging research (deng2025visual).
@@ 3 ImageSet2Text
ImageSet2Text generates textual descriptions of image sets that highlight the common visual elements present in most of the images. As shown in Fig. 3, it leverages external prebuilt components, including an LLM, a VQA model, a lexical graph, and a CVL model, to construct an intermediate concept graph, which then serves as the basis for generating the description. Given a set of $N$ images $\mathcal{D}=\{x_{1},\dots,x_{N}\}$, ImageSet2Text automatically generates a textual description $d$ that summarizes the visual elements in $\mathcal{D}$. This is achieved by constructing an intermediate concept graph represented as a list of verified hypotheses $\mathcal{G}_{c}=\{h_{\checkmark}^{1},...,h_{\checkmark}^{T}\}$, where each $h_{\checkmark}^{\tau}$ is a triplet $\langle s,p,o\rangle$ identifying a subject $s$, a predicate $p$, and an object $o$ that capture the key visual elements and their relations in $\mathcal{D}$. To build $\mathcal{G}_{c}$, ImageSet2Text follows an iterative process with $T$ iterations depicted in Fig. 3, with the following steps:
@@
    Initialization ($\tau=0$): It starts by defining the first subject $s\coloneq\text{`{image}'}$ and a list of three candidate predicates, $P\coloneq\{\text{`{content}'},\;\text{`{background}'},\;\text{`{style}'}\}$. The initial $\mathcal{G}_{c}^{0}$ contains only ‘image’ as root node.
@@
    Iterations ($\tau=1,...,T-1$), composed of two phases:
@@
      Guess what is in the set – A random subset of images $\mathcal{S}\subset\mathcal{D}$, with $|\mathcal{S}|=M\ll N$, is analyzed to hypothesize a full triplet $\langle s,p,o\rangle$ of elements present in $\mathcal{D}$, where $M$ is a predefined parameter.
@@
      Look and keep – The formulated hypothesis is verified on $\mathcal{D}$. If confirmed, it is used to update $\mathcal{G}_{c}^{\tau}$.
@@
    Termination ($\tau=T$): After convergence at $\tau=T$, a coherent textual description $d$ is generated from the final graph representation $\mathcal{G}_{c}=\mathcal{G}_{c}^{T}$.
@@
Figure 3: Overview of ImageSet2Text, considering an example set from the PairedImageSets datasets (visdiff). The figure shows how the different modules of the iterative process allow inferring information from the input image set, eventually generating a nuanced textual description.
@@ 3.1 Guess what is in the Set
The first phase generates a hypothesis from the random sample $\mathcal{S}$ of $M$ images for later verification on the full set $\mathcal{D}$. Let $\tau$ denote the current time step. From the current graph $\mathcal{G}_{c}^{\tau}$, ImageSet2Text selects the first encountered closest leaf node to the root node as the subject $s$, along with a random candidate predicate $p\in P$ appropriate for $s$. Fig. 3 depicts an illustrative example where the image set $\mathcal{D}$ contains desert photos. At iteration $\tau=5$, the subject selected is $s=\text{`{desert}'}$ and the predicate assigned is $p=\text{`{color}'}$.
@@
An LLM is prompted to ask a specific question about the images in $\mathcal{S}$ depending on the values of $s$ and $p$. In the given example, with $s=\text{`{desert}'}$ and $p=\text{`{color}'}$, the resulting question is:
@@
The question is posed to a VQA model for each of the $M$ images $x_{i}\in\mathcal{S}$, resulting in a set of answers $\mathcal{A}=\{a_{1},a_{2},\dots,a_{M}\}$, where $a_{i}$ denotes the answer for $x_{i}$.
@@
Given $\mathcal{A}$, each iteration verifies whether the predicate $p$ can expand $\mathcal{G}_{c}^{\tau}$. To achieve this, the LLM is prompted to perform two tasks: 1) summarization, i.e., condense $\mathcal{A}$ into a single hypothesis $h_{0}$, which is formulated as a triplet where the subject $s$ and predicate $p$ are given, and the object $o_{0}$ has to be derived from $\mathcal{A}$; and 2) completion, i.e., suggest possible expansions of $\mathcal{G}_{c}^{\tau}$ from $o_{0}$ as a list of new predicates to append to $P$. Following our example, $\mathcal{A}$ would yield hypothesis $h_{0}$ and continuations $P$:
$h_{0}=\langle\text{`{desert}'},\text{`{color}'},\text{`{gold}'}\rangle$ $P\leftarrow\{\text{`{shade}'},\text{`{brightness}'}\}$
@@
Since $h_{0}$ is derived from the subset $\mathcal{S}$, it may not generalize well to the full image set $\mathcal{D}$ due to sampling bias. To mitigate this, ImageSet2Text creates a set $\mathcal{H}=\{h_{0},h_{1},\dots,h_{k}\}$ of hypotheses ordered from the most general ($h_{k}$) to the most specific ($h_{0}$), where $h_{k}\supset h_{k-1}\supset\dots\supset h_{0}$. To generate $\mathcal{H}$ from $h_{0}$, ImageSet2Text relies on a given lexical graph $\mathcal{G}_{l}$. Let $\mathcal{G}_{l}=(V,R)$ be a directed graph, where $V$ is the set of lexical entries and $R$ represents semantic relations between nodes in $V$. For any given node $v\in V$ (e.g., ‘gold’), its parent node denotes a more general concept, or hypernym (e.g., ‘yellow’), and vice versa: ‘gold’ is a hyponym of ‘yellow’, i.e., a more specific lexical concept. In addition, two nodes $v_{1},v_{2}\in V$ are sibling nodes if they share the same parent node, but correspond to different lexical concepts (e.g., ‘gold’ and ‘gamboge’, both hyponyms of ‘yellow’).
The set $\mathcal{H}$ is obtained by traversing upward in the knowledge hierarchy of $\mathcal{G}_{l}$ by a maximum number of steps $\delta$. Given a hypothesis $h_{i}=\langle s,p,o_{i}\rangle$, its generalization $h_{i+1}$ is created as $h_{i+1}=\langle s,p,o_{i+1}\rangle$, where $o_{i}=\text{parent}^{i}(o_{0})$ and the parent function is the operation of moving to the hypernym of a lexical entry in $\mathcal{G}_{l}$. In the ongoing example, the hypotheses in $\mathcal{H}$ follow the hierarchy:
$h_{0}=\langle\text{`{desert}'},\text{`{color}'},\text{`{gold}'}\rangle$ $h_{1}=\langle\text{`{desert}'},\text{`{color}'},\text{`{yellow}'}\rangle$ $h_{2}=\langle\text{`{desert}'},\text{`{color}'},\text{`{chromatic color}'}\rangle$